# Equilibrium Investment with Random Risk Aversion: (Non-)uniqueness, Optimality, and Comparative Statics

### v2

@@ 3 Equilibrium Strategy
We first transform Condition (2.5) into an ordinary differential equation (ODE) for $v_{a}$. Differentiating $v_{a}(t)=\int_{t}^{T}|a(s)|^{2}ds$ with respect to $t$, we get from (2.5) that $v_{a}$ satisfies the following ODE:
@@
Conversely, assume that $v$ solves the ODE (3.1) and let $a(\cdot)=h(v(\cdot))\lambda(\cdot)$. Then $a$ satisfies (2.5).
The ODE (3.1) is an equation with separated variables. We can obtain its solutions in closed form.
@@
$H(y)\triangleq\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\int_{0}^{y}\left({\tilde{\mathbb{E}}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\left/{\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\right.\right)^{2}\mathrm{d}x,\quad y\in[0,\infty].$
@@
If $\Lambda(0)=0$, then $\lambda(\cdot)\equiv 0$ on $[0,T)$ and, in view of (2.5), there is a unique equilibrium $\pi\equiv\mathbf{0}$. This trivial case is excluded by making the following standing assumption.
@@
The next theorem provides, in closed form, the solutions to the ODE (3.1) and the equilibrium strategies in $\Pi_{d}$ .
@@
    If $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, the ODE (3.1) admits a one-parameter family of solutions, indexed by $T_{0}\in\mathcal{T}$:
@@
    $\mathcal{T}\triangleq\left\{T_{0}\in[0,T]\;\bigg|\;H(\infty)>\Lambda(0)-\Lambda(T_{0})\text{ and }\Lambda(t)>\Lambda(T_{0})\text{ for all }t\in[0,T_{0})\right\}.$
    Then the set of all equilibrium strategies in $\Pi_{d}$ is :
@@
Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$. In this case, $h(x)>0$ for all $x\geq 0$. Suppose that $v(\cdot)$ is a solution to the ODE (3.1). Separating variables in (3.1) yields
$-|\lambda(t)|^{2}=\frac{v’(t)}{h^{2}(v(t))},\quad t\in[0,T).$ (3.3)
@@
$-\Lambda(t)=\int_{t}^{T}\frac{v’(s)}{h^{2}(v(s))}\mathrm{d}s=\int_{v(t)}^{0}\frac{\mathrm{d}y}{h^{2}(y)}=-\int_{0}^{v(t)}\frac{\mathrm{d}y}{h^{2}(y)}=-H(v(t)),$
where the second equality follows from the change of variables $y=v(s)$. Thus, $H(v(t))=\Lambda(t)$ for all $t\in[0,T]$. As $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}>0$, it follows that $H(0)=0$ and $H(\cdot)\in{C}^{1}\left([0,+\infty)\right)$ is strictly increasing. Thus, the inverse function $H^{-1}$ is well-defined, continuous, and strictly increasing in $\left[0,H(\infty)\right)$. Thus, $H(\infty)>H(v(0))=\Lambda(0)$ and $v(t)=H^{-1}(\Lambda(t))$ for all $t\in[0,T]$, which yields the uniqueness of the solution. Conversely, if $H(\infty)>\Lambda(0)$, it is straightforward to verify that $v(\cdot)=H^{-1}(\Lambda(\cdot))$ indeed solves uniquely the ODE.
Assume that $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$. We can see that $H(0)=0$, $H\in C^{0}([0,+\infty])\cap C^{1}((0,+\infty))$, and $H$ is strictly increasing. Suppose that $v(\cdot)$ is a solution to the ODE (3.1). Let $T_{0}=\inf\{t\in[0,T]:v(t)=0\}$. Then $v(t)=0$ for any $t\in[T_{0},T]$ and $v(t)>0$ for any $t\in[0,T_{0})$. For any $t\in[0,T_{0})$, integrating (3.3) from $t$ to $T_{0}$ yields
@@
Assume that $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)=\infty$ for some $\varepsilon>0$. This implies that $\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\infty$ for any $y>0$. Clearly, $v\equiv 0$ is a solution (recall that $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}=0$). Suppose, on the contrary, that there exists another solution $v(\cdot)$ with $v(t)>0$ for some $t<T$. Let $t^{\prime}\triangleq\inf\{s\in(t,T]:v(s)=0\}$. By continuity, we have $v(t^{\prime})=0$ and $v(s)>0$ for any $s\in[t,t^{\prime})$. Integrating the ODE (3.3) from $t$ to $t^{\prime}$ yields
@@
which is impossible. Therefore, the ODE (3.1) admits the unique solution $v\equiv 0$. ∎
Assertion (2) of Theorem 3.2 characterizes the solutions to the ODE (3.1) via the parameter $T_{0}\in\mathcal{T}$. The following lemma provides an alternative explicit characterization of the set $\mathcal{T}$ via an auxiliary mapping.
@@
$\quad\varphi(\eta)\triangleq\min\{t\in[0,T]\mid\Lambda(t)=\eta\},\quad\eta\in[0,\Lambda(0)].$
Then we have $\mathcal{T}=\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$.
@@
Thus, $\mathcal{T}\subset\{\varphi(\eta)\mid\eta\in\mathcal{A}\}$. Conversely, for any $\eta\in\mathcal{A}$, the continuity of $\Lambda$ implies $\Lambda(\varphi(\eta))=\eta\in\mathcal{A}$. Moreover, for any $t\in[0,\varphi(\eta))$, we have $\Lambda(t)>\Lambda(\varphi(\eta))$ by the definition of $\varphi$. Therefore, $\varphi(\eta)\in\mathcal{T}$ and hence $\{\varphi(\eta)\mid\eta\in\mathcal{A}\}\subset\mathcal{T}$. ∎
@@
Because $\varphi$ is strictly decreasing, the sets $\mathcal{T}$ and $\mathcal{A}$ are in one-to-one correspondence. It is easily seen that $\mathcal{A}$ contains infinitely many elements. Consequently, under the conditions $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, we obtain infinitely many equilibria.
It should be noted that the existing studies, such as Desmettre2023 and liang2025short, focus mainly on the specific cases covered by Theorem 3.2(1), under the additional assumptions that $h(\cdot)$ is bounded and $\bm{R}$ is square-integrable. Our result covers any general random risk aversion $\bm{R}$. In the following two examples, the functions $h$ are unbounded and the equilibrium strategies are explicitly worked out.
@@
    with $H(\infty)=\theta^{2}$. If $\Lambda(0)<\theta^{2}$, then $v(t)=H^{-1}(\Lambda(t))=-\ln\left(1-\dfrac{\Lambda(t)}{\theta^{2}}\right)$. In this case, the uinque equilibrium strategy is explicitly given by
    $\bar{\pi}(t)=(\sigma^{\top})(t)^{-1}a(t)=(\sigma^{\top}(t))^{-1}h(v(t))\lambda(t)=\dfrac{(\sigma^{\top}(t))^{-1}\lambda(t)}{\sqrt{\theta^{2}-\Lambda(t)}},\quad t\in[0,T).$
@@
This example is provided for assertions (2) and (3) in Theorem 3.2. Let $\bm{R}_{\alpha}\sim G_{\alpha}$, where $\alpha\in(0,1)$ is a fixed parameter and $G_{\alpha}$ denotes the distribution function whose Laplace transform satisfies [^4] See Pollard1946 and Feller1971.
@@
Note that $\tilde{\mathbb{E}}[\bm{R}_{\alpha}]=\infty$ for $\alpha\in(0,1)$. Moreover, $h(x)=\frac{1}{\alpha}\left(\frac{x}{2}\right)^{1-\alpha}$, $x\geq 0$. Obviously, $h$ is unbounded. For simplicity, we assume that $d=1$ and $\sigma(\cdot)\equiv\sigma\neq 0$ and $\lambda(\cdot)\equiv\lambda\neq 0$ are constants in $\mathbb{R}$.
  • Let $\alpha\in(0.5,1)$. By direct calculation, the solutions to the ODE (3.1) take the form
@@
  • Let $\alpha\in(0,0.5]$. Direct calculation shows that the ODE (3.1) admits only the trivial solution $v\equiv 0$ and therefore $\pi=0$ is the unique equilibrium.
Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$ and $H(\infty)>\Lambda(0)$. Assertion (1) of Theorem 3.2 shows that the ODE (3.1) admits a unique solution $v$ and therefore there exists a unique function $a$ such that (2.5) holds for all $t\in[0,T)$. To emphasize the dependence on $\bm{R}$, we occasionally write $v_{\bm{R}}$ and $a_{\bm{R}}$ for $v$ and $a$, respectively.
@@
The explicit characterization of the equilibrium strategy in Theorem 3.2(1) yields a clear structural decomposition of the risk exposure vector, given by
@@
Note that the definitions of the functions $h(\cdot)$ and $H(\cdot)$ depend solely on the distribution of the random risk aversion $\bm{R}$}. The representation (3.4) shows that, at time $t$, the effects of three distinct factors on the risk exposure vector are cleanly separated: the instantaneous market opportunity $\lambda(t)$, the aggregate future market opportunities $\Lambda(t)$ over the remaining interval $(t,T)$, and the random risk aversion preference itself.
More specifically, at time $t$, if $\Lambda(t)$ and $\bm{R}$} are fixed, a larger current risk premium $|\lambda(t)|$ leads to a larger $|a(t)|$. If $|\lambda(t)|$ and $\bm{R}$} are fixed, then a larger future market opportunity $\Lambda(t)$ implies a larger value of $H^{-1}(\Lambda(t))$, and hence a larger value of $h\big(H^{-1}(\Lambda(t))\big)$ (see Proposition 3.8 for a detailed discussion; see alos Lemma A.1 for the monotonicity of $h$). Finally, if both $\lambda(t)$ and $\Lambda(t)$ are fixed, one can study how $\bm{R}$} affects $h$ and $H$, and thereby its impact on $a(t)$; see the comparative statics analysis in Section 5.
We emphasize that, in contrast to the classical CRRA utility, which generates a “myopic” strategy reacting primarily to the instantaneous market price of risk $\lambda(t)$, the equilibrium under random risk aversion is inherently forward-looking due to the dependence on $\Lambda(t)$. Finally, we refer to liang2024dynamic and liang2025dynamic for similar structural representations, although monotonicity properties and comparative statics are not investigated therein.
Next, we study the impact of changes in future market opportunities on the risk exposure vector.
@@
Let $\bm{R}$ be a non-negative random variable with $\tilde{\mathbb{E}}[\bm{R}]\in(0,+\infty)$. Fix a vector $\lambda_{0}\in\mathbb{R}^{d}$. Consider a sequence of market settings indexed by $n\in\mathbb{N}$, defined by the pairs $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$, where $T_{n}>0$ and $\lambda_{n}:[0,T_{n}]\to\mathbb{R}^{d}$ satisfying $\lambda_{n}(0)=\lambda_{0}$. Let $\Lambda_{n}\triangleq\int_{0}^{T_{n}}|\lambda_{n}(s)|^{2}ds$. Let $a_{n}(0)$ denote the risk exposure vector at time 0 corresponding to the $n$-th market setting.
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$, then
    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}}\quad\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(if $r_{0}>0$)},$
    where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|$ converges monotonically increasing, i.e.,
    $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}.$
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$, then
    $\lim_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{\tilde{\mathbb{E}}[\bm{R}]}.$
    Moreover, $|a_{n}(0)|$ converges monotonically decreasing, i.e.,
    $|a_{n}(0)|\searrow\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}.$
Here, we adopt the convention that $\frac{1}{0}=\infty$.
@@
The convergence results in Proposition 3.8 depend solely on the sequence of integral values $\{\Lambda_{n}\}_{n\geq 1}$, and no convergence of the parameter sequence $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$ itself is required. Moreover, focusing on time 0 entails no loss of generality. For any arbitrary time $t$, analogous conclusions follow by a simple time translation, with $t$ treated as the initial time.
@@
In the single-stock case, assuming that $\lambda>0$ and $\sigma>0$ are constants, one can use $\hat{\pi}(\cdot)=\frac{1}{\sigma}h(v(\cdot))\lambda$, together with the monotonicity of $h$ given in Lemma A.1, to establish a favorable property: $\pi(\cdot)$ is decreasing. See Section 4.3 of Desmettre2023 for a detailed discussion. However, the results in Desmettre2023 do not yield Proposition 3.8. The proof of Proposition 3.8 crucially relies on our explicit characterization of the risk exposure vector, given in (3.4), which is not available in their framework.
As the final part of this section, we use the explicit characterization of the equilibria in Theorem 3.2 to investigate the continuity properties of $a(\cdot)$ with respect to the distribution of the random risk aversion. More specifically, the following theorem shows that if a sequence of random risk aversion $\{R_{n}\}_{n\geq 1}$ converges in distribution to $R$, then the corresponding sequence $\{a_{n}\}_{n\geq 1}$ of volatility converges to $a$.
@@
Let $\{\bm{R}_{n}\}_{n\geq 1}$ be a sequence of random risk aversions that converges in distribution to $\bm{R}$ and $\lim\limits_{n\to\infty}\tilde{\mathbb{E}}[\bm{R}_{n}]=\tilde{\mathbb{E}}[\bm{R}]\in(0,\infty)$. Assume that $H_{n}(\infty)>\Lambda(0)$ for all $n\geq 1$ and $H(\infty)>\Lambda(0)$. Let $a_{n}=a_{\bm{R}_{n}}$ and $a=a_{\bm{R}}$. Then $\lim\limits_{n\to\infty}a_{n}(t)=a(t)$ for all $t\in[0,T)$. Moreover, if $\lambda\in L^{\infty}([0,T])$, then the convergence is uniform on $[0,T)$.

### v3

@@ 3 Equilibrium Strategy
We first transform Condition (2.5) into an ordinary differential equation (ODE) for $v_{a}$. Differentiating $v_{a}(t)=\int_{t}^{T}|a(s)|^{2}\mathrm{d}s$ w.r.t. $t$, we get from (2.5) that $v_{a}$ satisfies the following ODE:
@@
Conversely, assume that $v$ solves ODE (3.1) and let $a(\cdot)\triangleq h(v(\cdot))\lambda(\cdot)$. Then $a$ satisfies (2.5).
ODE (3.1) is an equation with separated variables. We can obtain its solutions in closed form.
@@
$\displaystyle H(y)\triangleq\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\int_{0}^{y}\left({\tilde{\mathbb{E}}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\left/{\tilde{\mathbb{E}}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x}\right]}\right.\right)^{2}\mathrm{d}x,\quad y\in[0,\infty].$ (3.2)
@@
If $\Lambda(0)=0$, then $\lambda(\cdot)\equiv\mathbf{0}$ on $[0,T)$ and, in view of (2.5), there is a unique equilibrium $\bar{\pi}\equiv\mathbf{0}$. This trivial case is excluded by making the following standing assumption.
@@
The next theorem provides, in closed form, the solutions to ODE (3.1) and the equilibrium strategies in $\Pi_{d}$. We will use $\Pi_{d}^{e}$ to denote all equilibrium strategies in $\Pi_{d}$, that is,
$\Pi_{d}^{e}\triangleq\{\pi\in\Pi_{d}\,:\ \pi\text{ is an equilibrium strategy}\}.$
@@
    If $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, then ODE (3.1) admits a one-parameter family of solutions, indexed by $T_{0}\in\mathcal{T}$:
@@
    $\mathcal{T}\triangleq\left\{T_{0}\in[0,T]:H(\infty)>\Lambda(0)-\Lambda(T_{0})\text{ and }\Lambda(t)>\Lambda(T_{0})\text{ for all }t\in[0,T_{0})\right\}.$
    In this case,
@@
Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$. In this case, $h(x)>0$ for all $x\geq 0$. Suppose that $v(\cdot)$ is a solution to ODE (3.1). Separating variables in (3.1) yields
$-|\lambda(t)|^{2}=\frac{v^{\prime}(t)}{h^{2}(v(t))},\quad t\in[0,T).$ (3.4)
@@
$-\Lambda(t)=\int_{t}^{T}\frac{v^{\prime}(s)}{h^{2}(v(s))}\mathrm{d}s=\int_{v(t)}^{0}\frac{\mathrm{d}y}{h^{2}(y)}=-\int_{0}^{v(t)}\frac{\mathrm{d}y}{h^{2}(y)}=-H(v(t)),$
where the second equality follows from the change of variables $y=v(s)$. Thus, $H(v(t))=\Lambda(t)$ for all $t\in[0,T]$. Because $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}>0$, it follows that $H(0)=0$ and $H(\cdot)\in{C}^{1}\left([0,\infty)\right)$ is strictly increasing. Thus, the inverse function $H^{-1}$ is well-defined, continuous, and strictly increasing in $\left[0,H(\infty)\right)$. Thus, $H(\infty)>H(v(0))=\Lambda(0)$ and $v(t)=H^{-1}(\Lambda(t))$ for all $t\in[0,T]$, which yields the uniqueness of the solution. Conversely, if $H(\infty)>\Lambda(0)$, it is straightforward to verify that $v(\cdot)\triangleq H^{-1}(\Lambda(\cdot))$ indeed solves uniquely the ODE.
Assume that $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$. We can see that $H(0)=0$, $H\in C^{0}([0,\infty])\cap C^{1}((0,\infty))$, and $H$ is strictly increasing. Suppose that $v(\cdot)$ is a solution to ODE (3.1). Let $T_{0}\triangleq\inf\{t\in[0,T]:\ v(t)=0\}$. Then $v(t)=0$ for any $t\in[T_{0},T]$ and $v(t)>0$ for any $t\in[0,T_{0})$. For any $t\in[0,T_{0})$, integrating (3.4) from $t$ to $T_{0}$ yields
@@
Assume that $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)=\infty$ for some $\varepsilon>0$. This implies that $\int_{0}^{y}\frac{1}{h^{2}(x)}\mathrm{d}x=\infty$ for any $y>0$. Clearly, $v\equiv 0$ is a solution (recall that $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}=0$). Suppose, on the contrary, that there exists another solution $v(\cdot)$ with $v(t)>0$ for some $t<T$. Let $t^{\prime}\triangleq\inf\{s\in(t,T]:\ v(s)=0\}$. By continuity, we have $v(t^{\prime})=0$ and $v(s)>0$ for any $s\in[t,t^{\prime})$. Integrating ODE (3.4) from $t$ to $t^{\prime}$ yields
@@
which is impossible. Therefore, ODE (3.1) admits the unique solution $v\equiv 0$. ∎
Assertion (2) of Theorem 3.2 provides, in closed form, the solutions to ODE (3.1) via the parameter $T_{0}\in\mathcal{T}$. The following lemma provides an alternative explicit characterization of the set $\mathcal{T}$ via an auxiliary mapping.
@@
$\quad\varphi(\eta)\triangleq\min\{t\in[0,T]:\,\Lambda(t)=\eta\},\quad\eta\in[0,\Lambda(0)].$
Then we have $\mathcal{T}=\{\varphi(\eta):\,\eta\in\mathcal{A}\}$.
@@
Thus, $\mathcal{T}\subset\{\varphi(\eta):\,\eta\in\mathcal{A}\}$. Conversely, for any $\eta\in\mathcal{A}$, the continuity of $\Lambda$ implies $\Lambda(\varphi(\eta))=\eta\in\mathcal{A}$. Moreover, for any $t\in[0,\varphi(\eta))$, we have $\Lambda(t)>\Lambda(\varphi(\eta))$ by the definition of $\varphi$. Therefore, $\varphi(\eta)\in\mathcal{T}$ and hence $\{\varphi(\eta):\,\eta\in\mathcal{A}\}\subset\mathcal{T}$. ∎
@@
The sets $\mathcal{T}$ and $\mathcal{A}$ are in one-to-one correspondence because $\varphi$ is strictly decreasing. It is easily seen that $\mathcal{A}$ contains infinitely many elements. Consequently, under the conditions $\tilde{\mathbb{E}}[\bm{R}]=\infty$ and $H(\varepsilon)<\infty$ for all $\varepsilon>0$, we obtain infinitely many equilibria.
It should be noted that the existing studies, such as Desmettre and Steffensen (2023) and Liang, Wang, and Xia (2025a), focus mainly on the specific cases covered by Theorem 3.2(1), under the additional assumptions that $h(\cdot)$ is bounded and $\bm{R}$ is square-integrable. Our result covers any general RRA $\bm{R}$. In the following two examples, the functions $h$ are unbounded and the equilibrium strategies are explicitly worked out.
@@
    with $H(\infty)=\theta^{2}$. If $\Lambda(0)<\theta^{2}$, then $v(t)=H^{-1}(\Lambda(t))=-\ln\left(1-\dfrac{\Lambda(t)}{\theta^{2}}\right)$. In this case, the unique equilibrium strategy is explicitly given by
    $\bar{\pi}(t)=(\sigma^{\top}(t))^{-1}a(t)=(\sigma^{\top}(t))^{-1}h(v(t))\lambda(t)=\dfrac{(\sigma^{\top}(t))^{-1}\lambda(t)}{\sqrt{\theta^{2}-\Lambda(t)}},\quad t\in[0,T).$
@@
This example is provided for Assertions (2) and (3) in Theorem 3.2. Let $\bm{R}_{\alpha}\sim G_{\alpha}$, where $\alpha\in(0,1)$ is a fixed parameter and $G_{\alpha}$ denotes the distribution function whose Laplace transform satisfies (see Pollard (1946) and Feller (1971, Section XIII.6))
@@
Note that $\tilde{\mathbb{E}}[\bm{R}_{\alpha}]=\infty$ for $\alpha\in(0,1)$. Moreover, $h(x)=\frac{1}{\alpha}\left(\frac{x}{2}\right)^{1-\alpha}$, $x\geq 0$. Obviously, $h$ is unbounded. For simplicity, we assume that $d=1$, $\sigma(\cdot)\equiv\sigma\neq 0$ and $\lambda(\cdot)\equiv\lambda\neq 0$ are constants in $\mathbb{R}$.
  • Let $\alpha\in(0.5,1)$. By direct calculation, the solutions to ODE (3.1) take the form
@@
  • Let $\alpha\in(0,0.5]$. Direct calculation shows that ODE (3.1) admits only the trivial solution $v\equiv 0$ and therefore $\bar{\pi}=0$ is the unique equilibrium.
Assume $\tilde{\mathbb{E}}[\bm{R}]<\infty$ and $H(\infty)>\Lambda(0)$. Assertion (1) of Theorem 3.2 shows that ODE (3.1) admits a unique solution $v$ and therefore the deterministic equilibrium strategy exists and is unique. To emphasize the dependence on $\bm{R}$, we occasionally write $v_{\bm{R}}$ and $a_{\bm{R}}$ for $v$ and $a$, respectively.
@@
The closed form of equilibrium strategy in Theorem 3.2(1) yields a clear structural decomposition of the risk exposure vector $a$, given by
@@
Note that the functions $h(\cdot)$ and $H(\cdot)$ depend solely on the distribution of the RRA $\bm{R}$}. The representation (3.5) shows that, at time $t$, the effects of three distinct factors on the risk exposure vector are totally separated: the instantaneous market opportunity $\lambda(t)$, the aggregate future market opportunities $\Lambda(t)$ over the remaining interval $(t,T)$, and the RRA itself (represented by $h$ and $H$).
More specifically, at time $t$, if $\Lambda(t)$ and $\bm{R}$} are fixed, a larger current risk premium $|\lambda(t)|$ leads to a larger $|a(t)|$. If $|\lambda(t)|$ and $\bm{R}$} are fixed, then a larger future market opportunity $\Lambda(t)$ implies a larger value of $H^{-1}(\Lambda(t))$, and hence a larger value of $h\big(H^{-1}(\Lambda(t))\big)$ (see Proposition 3.8 for a detailed discussion; see also Lemma A.1 for the monotonicity of $h$). Finally, if both $\lambda(t)$ and $\Lambda(t)$ are fixed, one can study how $\bm{R}$} affects $h$ and $H$, and thereby its impact on $a(t)$; see the comparative statics analysis in Section 5.
We emphasize that, in contrast to the classical CRRA utility, which generates a “myopic” strategy reacting primarily to the instantaneous market price of risk $\lambda(t)$, the equilibrium under RRA is inherently forward-looking due to the dependence on $\Lambda(t)$. Finally, we refer to Liang et al. (2024) and Liang, Xia, and Yuan (2025b) for similar structures of the equilibrium strategies under other preferences, although monotonicity properties and comparative statics are not investigated therein.
Next, we study the effect of changes in future market opportunities on the risk exposure vector.
@@
Let $\bm{R}$ be a non-negative random variable with $\tilde{\mathbb{E}}[\bm{R}]\in(0,\infty)$. Fix a vector $\lambda_{0}\in\mathbb{R}^{d}$. Consider a sequence of market settings indexed by $n\in\mathbb{N}$, defined by the pairs $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$, where $T_{n}>0$ and $\lambda_{n}:[0,T_{n}]\to\mathbb{R}^{d}$ satisfy $\lambda_{n}(0)=\lambda_{0}$. Let $\Lambda_{n}\triangleq\int_{0}^{T_{n}}|\lambda_{n}(s)|^{2}ds$. Let $a_{n}(0)$ denote the risk exposure vector at time 0 corresponding to the $n$-th market setting.
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly increases to $H(\infty)$, then $\lim\limits_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{r_{0}}$, where $r_{0}\triangleq\operatorname{essinf}\bm{R}$. Moreover, $|a_{n}(0)|\nearrow\frac{|\lambda_{0}|}{r_{0}}$ as $n\to\infty$.
@@
    If the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly decreases to 0, then $\lim\limits_{n\to\infty}a_{n}(0)=\frac{\lambda_{0}}{\tilde{\mathbb{E}}[\bm{R}]}$. Moreover, $|a_{n}(0)|\searrow\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}$ as $n\to\infty$.
Here, we adopt the conventions $\frac{1}{0}=\infty$ and $0\cdot\infty=0$.
@@
The convergence results in Proposition 3.8 depend solely on the sequence of integral values $\{\Lambda_{n}\}_{n\geq 1}$, and no convergence of the parameter sequence $\{(\lambda_{n}(\cdot),T_{n})\}_{n\geq 1}$ itself is required. Moreover, focusing on time 0 does not cause a loss of generality. For any time $t$, analogous conclusions follow by a simple time translation, with $t$ treated as the initial time.
@@
In the single-stock case, assuming that $\lambda>0$ and $\sigma>0$ are constants, one can use $\widehat{\pi}(\cdot)=\frac{1}{\sigma}h(v(\cdot))\lambda$, together with the monotonicity of $h$ given in Lemma A.1, to establish a favorable property: $\pi(\cdot)$ is decreasing. See Section 4.3 of Desmettre and Steffensen (2023) for a detailed discussion. However, the results in Desmettre and Steffensen (2023) do not yield Proposition 3.8. The proof of Proposition 3.8 is crucially based on our closed-form expression for the risk exposure vector, given in (3.5), which is not available in their work.
As the final part of this section, we use the closed form of the equilibria in Theorem 3.2 to investigate the continuity properties of $a(\cdot)$ w.r.t. the distribution of the random risk aversion. More specifically, the following theorem shows that if a sequence of RRA $\{\bm{R}_{n}\}_{n\geq 1}$ converges in distribution to $\bm{R}$, then the corresponding sequence of risk exposure vectors $\{a_{n}\}_{n\geq 1}$ converges to $a$.
@@
Let $\{\bm{R}_{n}\}_{n\geq 1}$ be a sequence of random risk aversions that converges in distribution to $\bm{R}$ and $\lim\limits_{n\to\infty}\tilde{\mathbb{E}}[\bm{R}_{n}]=\tilde{\mathbb{E}}[\bm{R}]\in(0,\infty)$. Assume that $H_{n}(\infty)>\Lambda(0)$ for all $n\geq 1$ and $H(\infty)>\Lambda(0)$. Let $a_{n}=a_{\bm{R}_{n}}$ and $a=a_{\bm{R}}$. Then $\lim\limits_{n\to\infty}a_{n}(t)=a(t)$ for all $t\in[0,T)$. Moreover, if $\lambda\in L^{\infty}([0,T))$, then the convergence is uniform on $[0,T)$.